# JUICER: Data-Efficient Imitation Learning for Robotic Assembly

### v2

@@ II-A Notation and System Components
We aim to learn a policy mapping observations to actions using a dataset of trajectories, , , with being the trajectory length. Our observations include RGB images (front and wrist views) and the robot’s proprioceptive state. Our actions represent desired 6-DoF end-effector poses, reached via a low-level Operation Space Controller [21]. We train with Behavior Cloning (BC), i.e., we optimize to maximize the likelihood of the data, .
@@ II-B Denoising Diffusion Probabilistic Models
Our method uses Denoising Diffusion Probabilistic Models (DDPM) [22, 23, 24] to learn high-dimensional distributions. The data-generating process is modeled with a fixed “forward” noising process and a learned reverse process . A sample is denoted , and are latent variables in the generation process ending as pure Gaussian noise with an appropriate choice of noise schedule . Using , we can sample any step as , allowing the reverse process to be estimated through a “noise prediction” surrogate loss[22],
The predicted noise is represented with a neural network. After training has converged with the above objective, we can sample from the distribution through an iterative “refinement” process, starting from Gaussian noise , and produce by subtracting the predicted noise repeatedly:
@@ III-B Data Collection and Annotation
The first step is robot teleoperation. While any teleoperation interface can be used, our system utilizes a 3DConnexion SpaceMouse for 6-DoF end-effector control. For each task, 50 demos are collected, taking 1-3 hours per task. Teleoperation in simulation offers unique opportunities to make demo collection more seamless and improve user experience. For instance, using the ability to perform resets, we implement an “undo” button and enable users to correct their mistakes while collecting trajectories. These additions save time by reducing the number of trials discarded due to operator failure.
After data collection, we add sparse annotations of “bottleneck” state transitions requiring high precision (these labels are used for data augmentation, see Section III-C) To annotate the demonstrations, a user loads the successful trajectories, steps through each episode, and marks the timesteps when bottleneck states occur. This step adds 15% to the total data collection time, and even the longer demos (2500 time steps) can be fully annotated in 1-2 minutes.
@@ III-C Trajectory Augmentation
Our experiments showed that small misalignment errors during precise grasps or insertions could quickly compound and lead to catastrophic task failure (e.g., parts falling out of reach in extreme out-of-distribution poses). We, therefore, want to improve the model’s ability to reach and transition through these bottleneck regions robustly. However, we want to avoid having the demonstrator manually provide examples of additional recovery behaviors [27, 28]. Our solution is to increase dataset support in regions that require high precision via synthetic data augmentation (see Algorithm 1). The first step is to reset the simulated robot and environment to a recorded bottleneck state from a demonstration. We then sample a random end-effector pose away from the bottleneck state and construct a sequence of “disassembly” actions that reach this pose via linear interpolation/SLERP. Finally, starting from the random pose, the robot executes a “correction” that returns from the randomly perturbed state to the original bottleneck. This correction is obtained by reversing the disassembly actions. We record the corrective actions and rendered scene images and compare the resulting scene state to the original demonstration. If the final states match, the trajectory segment is saved in the same format as the demonstrations to be used for BC training.
@@ III-D1 Data Processing and Image Encoding
We normalize all actions and proprioceptive states so each dimension lies in the range using and statistics from demo data across all tasks. Consistent with prior work, we have found that position control greatly outperforms delta control[6]. We represent orientations using the 6-dimensional representation[31, 32] resulting in an action space of . When processing image observations, we first resize raw RGB images from to . We then perform various random crops and color jitter to increase dataset diversity and reduce overfitting, and perform a final resizing down to .
We encode the processed images with a vision model. The encoder outputs latent image representations that we project to 128-dimensional vectors (one for each camera view). The final observation provided to the policy is of size . Contrary to [6], we find an ImageNet-pre-trained ResNet18 with global pooling to outperform a ResNet18 trained from scratch with SpatialSoftmax pooling, similar to[7]. Each ResNet18 has 11 million parameters.
@@ III-D2 Diffusion Policy
Given a dataset of observation-action chunks, , the goal is to fit a policy model . The standard denoising objective in fitting diffusion models from [22] is adapted to the setting of modeling an action distribution conditioned on an observation, , but where the noise model is parameterized as . Both the action and observation are chunks along the time-dimension, and . We find longer prediction horizons benefit performance and use in all our experiments. We still find the action execution horizon of provides a good tradeoff between open-loop consistency and closed-loop reactivity. Consistent with recent work[20], we have not found including history helpful and use for all tasks.
Like prior work[33, 6], we use a 1D temporal CNN-based U-Net introduced by[4], with the same FiLM conditioning layers in the residual blocks used in [6]. Experiments with the time-series diffusion transformer introduced in[6] did not improve performance in our tasks. Unless stated otherwise, our experiments use a U-Net with 3 downsampling blocks of dimension 256, 512, and 1024, yielding a model with 69 million parameters and a full policy model with 91 million parameters. At training time, we used 100 denoising steps for the DDPM sampler[22] and 8 denoising steps with the DDIM sampler[25, 26] at inference time.
@@ IV-A1 Tasks
We choose 4 tasks from the benchmark FurnitureBench[1]: round_table, lamp, square_table, and one_leg (where one_leg is a subtask of the square_table, see Figure 3). This task subset presents a variety of challenges which are summarized in Table I.
@@ IV-A3 Evaluation
We evaluate the 5 models trained from different random initializations per condition on 100 rollouts and report the mean and max success rate, i.e., the share of attempts that result in a complete assembly. We also examine what performance can be achieved using an even smaller budget of human-collected demos (see Section IV-B7).
@@ IV-A4 Baselines
Figure 4: Average and maximum success rates (%) of methods across tasks. Bolded methods are components of our JUICER pipeline.
Figure 5: Learning one_leg from 10 human demos.
@@ IV-B7 Further Minimizing Demonstration Budget
Finally, we perform one full JUICER iteration starting from just 10 demonstrations of the one_leg task and compare it with the performance obtained from 50 human-collected demonstrations. Figure 5 shows that training the baseline diffusion policy on only 10 demonstrations for the one_leg task yields a success rate of 19%, compared to 59% with 50 demonstrations. However, models trained on a combined set of 10 human demonstrations, 90 successful rollouts, and 100 augmentations achieve, on average, 71% success, a improvement beyond collecting 50 demos by hand.
@@ Conclusion
ACKNOWLEDGMENT
The computations in this paper were run on the FASRC cluster, supported by the FAS Division of Science Research Computing Group at Harvard University, and on the MIT Supercloud[58].
@@ VI-A1 Training Hyperparameters
Parameter |
Value |
Control Mode |
Position |
Action Space Dimension |
10 |
Observation Space Dimension |
16 |
Orientation Representation |
6D |
Max LR |
LR Scheduler |
Cosine |
Weight Decay |
Warmup steps |
500 |
Batch Size |
256 |
Max Epochs |
400 |
Steps per Epoch |
400 |
Image Size Input |
Image Size Encoder |
Vision Encoder Model |
ResNet18 |
Encoder Weights |
ImageNet 1k |
Encoder Parameters |
million |
Runs per Condition |
5 |
Encoder Feature Projection Dim |
128 |
Parameter |
Value |
U-Net Down dims |
[256, 512, 1024] |
U-Net Parameters |
69 million |
Policy total parameters |
91 million |
Observation Horizon |
1 |
Prediction Horizon |
32 |
Action Horizon |
8 |
DDPM Training Steps |
100 |
DDIM Inference Steps |
8 |
Parameter |
Value |
Residual Blocks |
5 |
Residual Block Width |
1024 |
Parameters |
10 million |
Policy total parameters |
32 million |
Observation Horizon |
1 |
Prediction Horizon (Chunked) |
32 |
Action Horizon (Chunked) |
8 |
Prediction Horizon (No Chunking) |
1 |
Action Horizon (No Chunking) |
1 |
@@ VI-B3 Dataset Sizes Across Tasks and Methods
VI-C Results in Tabular Form
@@
VI-D Further Analysis
VI-D1 Data Efficiency
@@
VI-D2 Number of Denoising Steps in DDIM Sampler
@@
VI-D3 Ratio of Synthetic to Demonstration Data
@@
VI-D4 Importance of Action-Chunking
@@
VI-D5 Comparison of Different Vision Encoders
@@
VI-D6 Empirical Mode Sampling

### v3

@@ II-A Notation and System Components
We aim to learn a policy mapping observations to actions using a dataset of trajectories, , , with being the trajectory length. Our observations include RGB images (front and wrist views) and the robot’s proprioceptive state. The proprioceptive state contains the current end-effector pose and velocity, and current gripper width. Our actions represent desired 6-Degrees-of-Freedom (DoF) end-effector poses, reached via a low-level Operation Space Controller [21], and a gripper open/close action. We train with Behavior Cloning (BC), i.e., we optimize to maximize the likelihood of the data, .
@@ II-B Denoising Diffusion Probabilistic Models
Denoising Diffusion Probabilistic Models (DDPM) [22, 23, 24] are a class of generative models used to learn high-dimensional distributions. DDPMs work by gradually adding noise to data and then learning to reverse this process.
The model consists of two main processes: (1) A fixed forward process that progressively adds Gaussian noise to the data. (2) A learned reverse process that removes noise step-by-step to generate new samples.
During training, the model learns to predict and remove noise from partially noised inputs using a noise prediction loss.
At inference time, generation starts from pure Gaussian noise. The model then iteratively refines this noise into a coherent sample by repeatedly predicting and subtracting noise. Techniques like the DDIM sampler [25, 26] allows for faster, deterministic sampling with fewer steps.
For a detailed mathematical treatment of DDPMs, including the specific equations for the forward and reverse processes, please refer to Section VI-C.
@@ III-B Data Collection and Annotation
The first step is robot teleoperation. While any teleoperation interface can be used, our system utilizes a 3DConnexion SpaceMouse for 6-DoF end-effector control. For each task, 50 demos are collected, taking 1-3 hours per task. Teleoperation in simulation offers unique opportunities to make demo collection more seamless and improve user experience. For instance, using the ability to perform resets, we implement an “undo” button that restores the state 10 frames earlier, enabling users to correct mistakes instantly. These additions save time by reducing the number of trials discarded due to operator failure.
After data collection, we add sparse annotations of “bottleneck” state transitions requiring high precision (these labels are used for data augmentation, see Section III-C). To annotate the demonstrations, a user loads the successful trajectories, steps through each episode, and marks the timesteps when bottleneck states occur. This step adds 15% to the total data collection time, and even the longer demos (2500 time steps) can be fully annotated in 1-2 minutes. We perform the annotation in a separate step as we found it hard for the demonstrator to remember to label bottleneck states accurately while providing demos, thus reducing the cognitive load.
@@ III-C Trajectory Augmentation
Our experiments showed that small misalignment errors during precise grasps or insertions could quickly compound and lead to catastrophic task failure (e.g., parts falling out of reach in extreme out-of-distribution poses). We, therefore, want to improve the model’s ability to reach and transition through these bottleneck regions robustly. However, we want to avoid having the demonstrator manually provide examples of additional recovery behaviors [27, 28]. Our solution is to increase dataset support in regions that require high precision via synthetic data augmentation (see Algorithm 1). The first step is to reset the simulated robot and environment to a recorded bottleneck state from a demonstration. We then sample a random end-effector pose away from the bottleneck state and construct a sequence of “disassembly” actions that reach this pose via (spherical) linear interpolation. Finally, starting from the random pose, the robot executes a “correction” that returns from the randomly perturbed state to the original bottleneck. This correction is obtained by reversing the disassembly actions. We record the corrective actions and rendered scene images and compare the resulting scene state to the original demonstration. If the final states match (within an error ), the trajectory segment is saved in the same format as the demonstrations to be used for BC training. Some sampled trajectories are infeasible causing other parts to move in the backward portion, necessitating the need for success check after the forward portion. In our experiments, augmentation transitions will constitute 10-20% of the total number of transitions (see Section VI-B3 for exact data mixture numbers).
@@ III-D1 Data Processing and Image Encoding
We normalize all actions and proprioceptive states so each dimension lies in the range using and statistics from demo data across all tasks. Consistent with prior work, we found position control to outperform delta control[6]. In position control, the action directly specifies the target position of the end-effector, while in delta control, the action represents the change in position relative to the current state. We represent orientations using a 6-dimensional representation shown to be easier to learn [31, 32], resulting in an action space of (see details in Section VI-A3). When processing image observations, we first resize raw RGB images from to . We then perform random cropping and color jitter to increase dataset diversity and reduce overfitting, and perform a final resizing to .
We encode the images with a vision model. The encoder outputs latent image representations that we project to 128-dimensional vectors (one for each camera view). The final observation provided to the policy is of size 11116 dimensions comes from 3 position coordinates, 6 for orientation, 3 for linear velocity, 3 for angular velocity, and 1 for gripper action.. Contrary to [6], we find an ImageNet-pre-trained ResNet18 with global pooling to outperform a ResNet18 trained from scratch with SpatialSoftmax pooling, similar to[7]. Each ResNet18 has 11M parameters.
@@ III-D2 Diffusion Policy
Given a dataset of observation-action chunks, , the goal is to fit a policy model . The standard denoising objective in fitting diffusion models from [22] is adapted to the setting of modeling an action distribution conditioned on an observation, , but where the noise model is parameterized as . Both the action and observation are chunks along the time-dimension, and . We find longer prediction horizons benefit performance and use in all our experiments. We still find the action execution horizon of provides a good tradeoff between open-loop consistency and closed-loop reactivity. Consistent with recent work[20], we have not found including history helpful and use for all tasks. Since velocity is included in the proprioception, multiple frames is not needed to infer it.
Like prior work[33, 6], we use a 1D temporal CNN-based U-Net introduced by[4], with the same FiLM conditioning layers in the residual blocks used in [6]. Experiments with the time-series diffusion transformer introduced in[6] did not improve performance in our tasks. Unless stated otherwise, our experiments use a U-Net with 3 downsampling blocks of dimension 256, 512, and 1024, yielding a model with 69 million parameters and a full policy model with 91 million parameters. At training time, we used 100 denoising steps for the DDPM sampler[22] and 8 denoising steps with the DDIM sampler[25, 26] at inference time. Hyperparameters were chosen by starting with best-practices presented in prior work as the nominal and then sweeping over different parameter choices. The full set of hyperparameters used is presented in Section VI-A1.
@@ IV-A1 Tasks
We choose 4 tasks from the benchmark FurnitureBench[1]: round_table, lamp, square_table, and one_leg (where one_leg is a subtask of the square_table, see Figure 3). We note that while the environment includes fiducial markers on the parts, our approach does not utilize these markers at all, relying solely on raw RGB images. This task subset presents a variety of challenges which are summarized in Table I.
@@ IV-A3 Evaluation
We evaluate the 5 models trained from different random initializations per condition on 100 rollouts and report the mean and max success rate, i.e., the share of attempts that result in a complete assembly. This is measured with comparing relative part poses to the desired pose given by the CAD model in the fully assembled configuration. We also examine what performance can be achieved using an even smaller budget of human-collected demos (see Section IV-B7).
@@ IV-B3 Trajectory Augmentation (TA)
We believe that TA provides larger performance gains on these tasks because the primary bottleneck in round_table and one_leg involves precise insertion. In contrast, the main challenge in lamp is to grasp a bulb that can roll on the table, introducing a quasi-dynamic element not present in the other tasks. Our current augmentation procedure is less suited to addressing this rolling behavior, as it focuses on static configurations around insertion points. For square_table, the task comprises a long sequence of sub-tasks, many of which are not insertions and thus currently not addressed by the augmentation. Consequently, focusing solely on bottleneck states creates a more delicate balance between increasing support for some phases and not diluting the importance of others.
@@ IV-B7 Further Minimizing Demonstration Budget
Finally, we perform one full JUICER iteration (i.e., collecting demonstrations, augmenting, and performing collect-infer) starting from just 10 demonstrations of the one_leg task and compare it with the performance obtained from 50 human-collected demonstrations. Figure 5 shows that training the baseline diffusion policy on only 10 demonstrations for the one_leg task yields a success rate of 19%, compared to 59% with 50 demonstrations. However, models trained on a combined set of 10 human demonstrations, 90 successful rollouts, and 100 augmentations achieve, on average, 71% success, a improvement beyond collecting 50 demos by hand.
@@ Conclusion
ACKNOWLEDGMENTS
This work was partly supported by the Sony Research Award and the US Government. The computations in this paper were run on the FASRC cluster, supported by the FAS Division of Science Research Computing Group at Harvard University, and on the MIT Supercloud[58].
Contributions LA led the project, implementation, and experimental evaluation, and contributed to ideation and writing. AS conceived the initial project idea, advised on project goals, contributed to the implementation of some parts of the system and writing. IS actively contributed in research discussions and provided feedback on the paper and writing. PA advised the project and contributed to its development, experimental design, and positioning.
@@ VI-A1 Training Hyperparameters
| Parameter | Value |
| Control Mode | Position |
| Action Space Dimension | 10 |
| Observation Space Dimension | 16 |
| Orientation Representation | 6D |
| Max LR | |
| LR Scheduler | Cosine |
| Weight Decay | |
| Warmup steps | 500 |
| Batch Size | 256 |
| Max Epochs | 400 |
| Steps per Epoch | 400 |
| Image Size Input | |
| Image Size Encoder | |
| Vision Encoder Model | ResNet18 |
| Encoder Weights | ImageNet 1k |
| Encoder Parameters | million |
| Runs per Condition | 5 |
| Encoder Feature Projection Dim | 128 |
| Parameter | Value |
| U-Net Down dims | [256, 512, 1024] |
| U-Net Parameters | 69 million |
| Policy total parameters | 91 million |
| Observation Horizon | 1 |
| Prediction Horizon | 32 |
| Action Horizon | 8 |
| DDPM Training Steps | 100 |
| DDIM Inference Steps | 8 |
| Parameter | Value |
| Residual Blocks | 5 |
| Residual Block Width | 1024 |
| Parameters | 10 million |
| Policy total parameters | 32 million |
| Observation Horizon | 1 |
| Prediction Horizon (Chunked) | 32 |
| Action Horizon (Chunked) | 8 |
| Prediction Horizon (No Chunking) | 1 |
| Action Horizon (No Chunking) | 1 |
@@ VI-B3 Dataset Sizes Across Tasks and Methods
VI-C Denoising Diffusion Probabilistic Models
Denoising Diffusion Probabilistic Models (DDPM) [22, 23, 24] are a powerful class of generative models used to learn and sample from high-dimensional distributions. This appendix provides a detailed explanation of the DDPM framework, including its mathematical foundations and key concepts.
Overview
DDPMs work on the principle of gradually adding noise to data and then learning to reverse this process. This approach allows the model to generate new samples by starting from pure noise and progressively refining it into coherent data.
The Forward Process
The forward process, also known as the diffusion process, is a fixed Markov chain that gradually adds Gaussian noise to the data. For a data point , we define a sequence of latent variables , where is the total number of steps in the process.
The forward process is defined as:
Here, is a noise schedule that determines how much noise is added at each step. The process is designed such that becomes pure Gaussian noise: .
The Reverse Process
The reverse process, which is learned during training, aims to gradually denoise the data, starting from pure noise. It is modeled as:
where represents the parameters of the model, typically a neural network.
Training Objective
To train the model, we use a “noise prediction” surrogate loss. This approach is made possible by the ability to sample from any intermediate step of the forward process directly:
where .
The training objective is then formulated as:
Here, is a neural network that predicts the noise added to the input.
Sampling Process
After training, we can generate new samples by starting from Gaussian noise and iteratively applying the learned reverse process:
where and is a small noise term added for stochasticity.
Improved Sampling
The DDIM sampler [25, 26] provides a method for deterministic sampling with fewer steps, improving the efficiency of the generation process while maintaining quality. This is achieved by modifying the noise prediction process and allowing for a more flexible trajectory between the noise and the final sample.
VI-D Results in Tabular Form
@@
VI-E Further Analysis
VI-E1 Data Efficiency
@@
VI-E2 Number of Denoising Steps in DDIM Sampler
@@
VI-E3 Ratio of Synthetic to Demonstration Data
@@
VI-E4 Importance of Action-Chunking
@@
VI-E5 Comparison of Different Vision Encoders
@@
VI-E6 Empirical Mode Sampling